# weighted-amenability and percolation

### v1

@@ 6.1. The question of $p_{h}<p_{u}$
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Then if $G$ is level-amenable then there cannot be infinitely many heavy clusters in Bernoulli$(p)$ percolation on $G$ for any $p\in[0,1]$.
A conjecture due to Hutchcroft [Hut20, Conjecture 8.5] states that this is, in fact, a characterization of the existence of the infinite phase for the heavy clusters. In light of Theorem 1.4, this conjecture could be restated as a generalization of the “$p_{c}<p_{u}$” conjecture of Benjamini and Schramm mentioned above, [BS96, Conjecture 6].
@@
Suppose that $p_{h}<p_{u}$. Then for any $p\in(p_{h},p_{u})$ there are infinitely many heavy clusters. By Theorem 5.1 for $\operatorname{\mathds{P}}_{p}$-a.e. configuration $\omega$, the random forest $\mathrm{FMaxSF}_{\mathbf{w}}(\omega)$ almost surely contains infinitely many trees with infinitely many $\mathbf{w}$-nonvanishing ends. Thus by Theorem 5.2 the cluster graphing with the measure induced by the law of such this forest is $\mu$-nonamenable. Finally, by Observation  2.23(c) the group $\Gamma$ also must be nonamenable. Theorem 2.3 concludes the proof. $\blacksquare$
Let us to mention that in [Tan19, Corollary 4.8] $p_{h}(G,\Gamma)\neq p_{u}(G)$ is characterized by the existence of an increasing exhausting sequence of slices $\{G_{n}\}_{n\in\mathbb{N}}$ such that $\lim_{n\to\infty}p_{c}(G_{n})<p_{u}(G).$
@@
Consider a $k>1$ and a graph $L_{k}$ with some unimodular group $\Delta$ of automorphisms acting on it in a quasi-transitive way with $k$ orbits. Let $V_{k-1}$ be an arbitrary union of $k-1$ of these orbits. We will prove that there is a graph $L_{k-1}$ on $V_{k-1}$, such that every element of $\Delta|_{V_{k-1}}$ defines an isomorphism of $L_{k-1}$, and such that $L_{k-1}$ is quasi-isometric to $L_{k}$. Moreover, the edges of $L_{k-1}$ will be defined as a local function of edges of $L_{k}$, (i.e., the $L_{k-1}$-edges on a vertex are given by a function from a suitable ball around the vertex, and the function is invariant under the rooted isomorphisms of this ball). Once we have this construction, the proof is completed as follows. First use Theorem 1.4 to find a nonamenable subgraph $L_{k}$ of $G$ that is induced by the union of $k$ levels for a suitable $k<\infty$. Remove the level of lowest weight from $L_{k}$, call the remaining vertices $V_{k-1}$, and define $L_{k-1}$ for $L_{k}$ as above. Repeat this to obtain $L_{k-2}$ from $L_{k-1}$, and so on until we arrive to an $L_{1}$ where the restriction of $\Gamma$ to $L_{1}$ is transitive, and $L_{1}$ is quasi-isometric to $L_{2}$ and (inductively) to $L_{k}$. In particular, $L_{1}$ in nonamenable because $L_{k}$ is. Now, to every level of $G$ add edges as in $L_{1}$. The resulting graph $G^{\prime}$ satisfies all the requirements of the theorem.
@@
$\frac{d(L^{(k)})}{d(G^{(k)})}\to 1\quad\textnormal{ and }\quad\frac{\Phi_{E}(L^{(k)})}{\Phi_{E}(G^{(k)})}\to 1\quad\textnormal{as }k\to\infty.$ (22)
By this and (21), if $k$ is large enough then $\frac{\Phi_{E}(L^{(k)})}{d(L^{(k)})}>\frac{1}{\sqrt{2}}$ and even $\frac{\Phi_{E}(L^{(k)})}{d(G^{(k)})}>\frac{1}{\sqrt{2}}$. We can furthermore choose $k$ such that $\frac{\Phi_{E}(G^{(k)})}{d(G^{(k)})}>\frac{1}{\sqrt{2}}$ (using (22)) and hence by [LP16, Theorem 7.38] we also have $\rho(G^{(k)})<\frac{1}{\sqrt{2}}$. Then we derive that
@@
where the second and the last inequalities are from Benjamini and Schramm [BS96], see Theorem 6.46 and respectively Theorem 7.32 and Lemma 7.33 in [LP16]. Choosing $G^{\prime}:=G^{(k)}$ completes the proof. $\blacksquare$
@@ 6.2. Continuity of the percolation phase transition for $\mathbf{w}$-nonamenable graphs
We now prove that there are no heavy clusters under critical Bernoulli$(p_{h})$ bond percolation on a $\mathbf{w}$-nonamenable graph $G$. In the unimodular setting it was shown in [BLPS99a]. The structure of our proof follows the same steps (see [LP16, Theorem 8.21]), our contribution is in the development of proper weighted analogs of the statements referenced in the argument.
@@
The inequality $p_{h}(G,\Gamma)<1$ immediately follows from Theorem 2.16.
The first part of the statement is proved by contradiction, which divides in two cases: if there is one or infinitely many heavy clusters.
Case of $1$. Suppose that there is a unique heavy cluster at $p_{h}$, denote it as $\sigma$. Such a cluster $\sigma$ is a $\Gamma$-invariant connected random subgraph of $G$ with $p_{h}(\sigma)=1$. Theorem 1.6 now yields that $G$ is $\mathbf{w}$-amenable giving a contradiction.
Case of $\infty$. Suppose there are infinitely many heavy clusters at $p_{h}$. By Theorem 5.1 the random spanning forest $\mathrm{FMaxSF}_{\mathbf{w}}(\omega_{p_{h}})$ contains a tree with infinitely many $\mathbf{w}$-nonvanishing ends almost surely. In particular, by Theorem 5.3 such a tree has $p_{h}<1$. Since $p_{h}(\omega_{p_{h}})=1$, we have that for a.e. $\mathrm{FMaxSF}_{\mathbf{w}}(\omega_{p_{h}})$ configuration $\varphi$ it also has $p_{h}(\varphi)=1$ yielding a contradiction. $\blacksquare$
@@ 7.2. Weighted random spanning trees and forests
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$. Is it true that $p_{h}(G,\Gamma)=p_{u}(G)$ if and only if $\mathrm{FMaxSF}_{\mathbf{w}}(G)$=$\mathrm{WMaxSF}_{\mathbf{w}}(G)$?
7.3. Tilted unimodular random graphs
In a follow-up work together with Gábor Elek, we introduce tilted unimodular random graphs as a generalization of nonunimodular quasi-transitive graphs, in the spirit of the unimodular random graphs of Aldous and Lyons [AL07]. Most arguments used in the present paper rely on the tilted mass transport principle, and could therefore be extended to this wider setting.
@@ Acknowledgments
We thank Russell Lyons, Anush Tserunyan, Robin Tucker-Drob, Konrad Wróbel and Tom Hutchcroft for helpful discussions, and Gabor Elek for suggesting Example 2.13. The first author was supported in part by the RTG award grant (DMS-2134107) from the NSF, the second author by the Icelandic Research Fund grant No. 239736-051. Both authors were also partially supported by the ERC Synergy Grant No. 810115 - DYNASNET.

### v2

@@ 6.1. The question of $p_{h}<p_{u}$
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Then if $G$ is level-amenable then there cannot be infinitely many heavy clusters in Bernoulli(p)(p) percolation on $G$ for any $p\in[0,1]$.
A conjecture due to Hutchcroft [Hut20, Conjecture 8.5] states that this is, in fact, a characterization of the existence of the infinite phase for the heavy clusters. In light of Theorem 1.4, this conjecture can be restated as a generalization of the $p_{c}<p_{u}$” conjecture of Benjamini and Schramm mentioned above, [BS96, Conjecture 6].
@@
Suppose that $p_{h}<p_{u}$. Then for any $p\in(p_{h},p_{u})$ there are infinitely many heavy clusters. By Theorem 5.1 for $\operatorname{\mathds{P}}_{p}$-a.e. configuration $\omega$, the random forest $\mathrm{FMaxSF}_{\mathbf{w}}(\omega)$ almost surely contains infinitely many trees with infinitely many $\mathbf{w}$-nonvanishing ends. Thus by Theorem 5.2 the cluster graphing of such a forest is $\mu$-nonamenable. Finally, by Observation  2.23(c) the group $\Gamma$ also must be nonamenable. Theorem 2.3 concludes the proof. $\blacksquare$
Let us also mention that in [Tan19, Corollary 4.8] $p_{h}(G,\Gamma)\neq p_{u}(G)$ is characterized by the existence of an increasing exhausting sequence of slices $\{G_{n}\}_{n\in\mathbb{N}}$ such that $\lim_{n\to\infty}p_{c}(G_{n})<p_{u}(G).$
@@
Consider a $k>1$ and a graph $L_{k}$ with some unimodular group $\Delta$ of automorphisms acting on it in a quasi-transitive way with $k$ orbits. Let $V_{k-1}$ be an arbitrary union of k1k-1 of these orbits. We will prove that there is a graph $L_{k-1}$ on $V_{k-1}$, such that every element of $\Delta|_{V_{k-1}}$ defines an isomorphism of $L_{k-1}$, and such that $L_{k-1}$ is quasi-isometric to $L_{k}$. Moreover, the edges of $L_{k-1}$ will be defined as a local function of edges of $L_{k}$, (i.e., the $L_{k-1}$-edges on a vertex are given by a function from a suitable ball around the vertex, and the function is invariant under the rooted isomorphisms of this ball). Once we have this construction, the proof is completed as follows. First use Theorem 1.4 to find a nonamenable subgraph $L_{k}$ of $G$ that is induced by the union of $k$ levels for a suitable $k<\infty$. Remove the level of lowest weight from $L_{k}$, call the remaining vertices $V_{k-1}$, and define $L_{k-1}$ for $L_{k}$ as above. Repeat this to obtain $L_{k-2}$ from $L_{k-1}$, and so on until we arrive to an $L_{1}$ where the restriction of $\Gamma$ to $L_{1}$ is transitive, and $L_{1}$ is quasi-isometric to $L_{2}$ and (inductively) to $L_{k}$. In particular, $L_{1}$ in nonamenable because $L_{k}$ is. Now, to every level of $G$ add edges as in $L_{1}$. The resulting graph $G^{\prime}$ satisfies all the requirements of the theorem.
@@
$\frac{d(L^{(k)})}{d(G^{(k)})}\to 1\quad\textnormal{as }k\to\infty.$ (23)
Consider any finite set $F\subseteq G^{(k)}$, and let $L_{1},L_{2},\ldots,L_{m}$ be the levels that it intersects. Then
$\frac{\left|\partial_{E(G^{(k)})}F\right|}{\left|F\right|}\geqslant\frac{\sum_{i=1}^{m}\left|\partial_{E(L_{i})}(F\cap L_{i})\right|}{\sum_{i=1}^{m}\left|F\cap L_{i}\right|}\geqslant\min_{i\in[m]}\frac{\left|\partial_{E(L_{i})}(F\cap L_{i})\right|}{\left|F\cap L_{i}\right|}\geqslant\Phi_{E}(L^{(k)}),$
where the first inequality holds as we discarded edges on the boundaries that connect different levels, and the last inequality holds because all levels are isomorphic to each other. Taking infimum over all $F$ yields that
$\Phi_{E}(G^{(k)})\geqslant\Phi_{E}(L^{(k)})\quad\textnormal{ for all }k\in\mathbb{N}.$ (24)
By this, (22) and (23), if $k$ is large enough then $\frac{\Phi_{E}(L^{(k)})}{d(L^{(k)})}>\frac{1}{\sqrt{2}}$ and even $\frac{\Phi_{E}(L^{(k)})}{d(G^{(k)})}>\frac{1}{\sqrt{2}}$. Thus, by (24), there exists $k$ such that $\frac{\Phi_{E}(G^{(k)})}{d(G^{(k)})}>\frac{1}{\sqrt{2}}$ and hence by [LP16, Theorem 7.38] we have $\rho(G^{(k)})<\frac{1}{\sqrt{2}}$. Then we derive that
@@
where the second and the last inequalities are from Benjamini and Schramm [BS96], see [LP16, Theorem 6.46] and [LP16, Theorem 7.32, Lemma 7.33] respectively. Choosing $G^{\prime}:=G^{(k)}$ completes the proof. $\blacksquare$
@@ 6.2. Continuity of the percolation phase transition for $\mathbf{w}$-nonamenable graphs
[…]
It is easy to show that in Theorem 1.6 one can replace (3) by “there is a $\Gamma$-invariant random connected hyperfinite subgraph of GG”. This modification leads to the following result.
Theorem 6.6.
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Suppose $G$ is $\mathbf{w}$-nonamenable. Suppose $p\in[0,1]$ is such that there is a heavy cluster $\operatorname{\mathds{P}}_{p}$-a.s. Then $\operatorname{\mathds{P}}_{p}$-a.s., every heavy cluster in $\omega$ is not hyperfinite, i.e., $\operatorname{\mathds{P}}_{p}(C(o)\textnormal{ is not hyperfinite}\mid\mathbf{w}(C(o))=\infty)=1.$
Remark 6.7.
Theorem 6.6 holds more generally for any insertion and deletion tolerant $\Gamma$-invariant percolation that admits a heavy cluster.
@@
In the case when there is a unique heavy cluster, one can extend a hyperfinite exhaustion of $\omega$ to an exhaustion of $G$ using the same argument as in Theorem 1.6(3).
In the case when there are infinitely many heavy clusters, Theorem 5.1 implies that a.s., for every heavy cluster $C\subseteq\omega$, there is a tree $T\subseteq\mathrm{FMaxSF}_{\mathbf{w}}(C)$ with $\geqslant 3$ $\mathbf{w}$-nonvanishing ends. So with positive probability $o\in T$ and on this event $T$ is not hyperfinite by Theorem 5.3. Hence $C$ is also not hyperfinite. $\blacksquare$
This immediately implies Theorem 1.8.
Proof of Theorem 1.8.
The first part of the statement follows from the fact that percolation configuration at $p_{h}$ is by definition hyperfinite, hence by Theorem 6.6 it contains only light clusters.
The inequality $p_{h}(G,\Gamma)<1$ follows from the first part or even more directly from Theorem 2.16. $\blacksquare$
We now give an alternative proof of one of the main results in [Tim06b], which implies [Tim06b, Corollary 5.8] (as in Theorem 6.1) used in the previous subsection. In fact, we present a slightly stronger (but conjecturally equivalent) statement as we do not require the existence of infinitely many heavy clusters. Instead, we assume that $G$ is $\mathbf{w}$-nonamenable.
Corollary 6.8 (Slight strengthening of [Tim06b, Theorem 5.5]).
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Suppose $G$ is $\mathbf{w}$-nonamenable and $p\in[0,1]$ is such that there is a heavy cluster $\operatorname{\mathds{P}}_{p}$-a.s. Then a.s. for any infinite cluster $C$ there exists a slice $S$ such that $C\cap S$ induces an infinite connected component.
Proof.
Suppose there is a heavy cluster $C\subset\omega$ such that for any slice $S$ we have that $C\cap S$ induces only finite components. Consider level-percolation of $G$ defined by a hyperfinite exhaustion of the level graph $\mathcal{L}(G)$. Since each connected component of such percolation is a slice, intersecting it with $\omega$ yields a hyperfinite exhaustion of $C$, contradicting Theorem 6.6. $\blacksquare$
Tracing down all of the statements on which this proof relies both in the present work and in [CTT22] one can see that it is indeed independent of [Tim06b]. However, some of the ideas in our argument are similar to those in [Tim06b]. For example, both approaches rely on partitioning of the levels, either by taking a 1-partition or a level-percolation. The proof in [Tim06b] was based on a construction of an infinitely-ended forest on a finite set of levels, thus reducing the problem to unimodular quasi-transitive graphs (similarly to the present proof of Theorem 1.7). Here, through the use of the $\mathrm{FMaxSF}$ with trees of infinitely many nonvanishing ends, we get a direct contradiction.
@@ 7.2. Weighted random spanning trees and forests
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$. Is it true that $p_{h}(G,\Gamma)=p_{u}(G)$ if and only if $\mathrm{FMaxSF}_{\mathbf{w}}(G)$$\mathrm{WMaxSF}_{\mathbf{w}}(G)$?
7.3. Modular random graphs
In a follow-up work together with Gábor Elek, we introduce (multi)modular random graphs as a generalization of nonunimodular quasi-transitive graphs, in the spirit of the unimodular random graphs of Aldous and Lyons [AL07]. Most arguments used in the present paper rely on the tilted mass transport principle, and could therefore be extended to this wider setting.
@@ Acknowledgments
We thank Russell Lyons, Robin Tucker-Drob, Konrad Wróbel and Tom Hutchcroft for helpful discussions, and Gabor Elek for suggesting Example 2.13. We also thank Anush Tserunyan for carefully reading the original draft and for many helpful suggestions that improved the article. The first author was supported in part by the RTG award grant (DMS-2134107) from the NSF, the second author by the Icelandic Research Fund grant No. 239736-051. Both authors were also partially supported by the ERC Synergy Grant No. 810115 - DYNASNET.